How many ounces of a 90% cedar oil solution need to be added to an empty gallon of water to make one gallon of a 10% cedar oil solution?
Answer by nerdybill(7384)

How many ounces of a 90% cedar oil solution need to be added to an empty gallon of water to make one gallon of a 10% cedar oil solution?
.
There are 128 ounces in a gallon
.
Let x = amount (ounces) of 90% cedar oil
then
128-x = amount (ounces) of water
.
.90(x) + 0(128-x) = .10(128)
.90(x) = 12.8
x = 12.8/.90
x = 14.22 ounces of cedar oil (answer)
.
amount of water:
128-x = 128-14.22 = 113.78 ounces of water